# Large Language Models Can Learn Temporal Reasoning

### v3

@@ Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal expressions and intricate temporal logic. In this paper, we propose TG-LLM, a novel framework towards language-based TR. Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that facilitates the TR learning. A synthetic dataset (TGQA), which is fully controllable and requires minimal supervision, is constructed for fine-tuning LLMs on this text-to-TG translation task. We confirmed in experiments that the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain of Thought (CoT) bootstrapping and graph data augmentation. We observed that those strategies, which maintain a balance between usefulness and diversity, bring more reliable CoTs and final results than the vanilla CoT distillation.  [^1] The code is available at https://github.com/xiongsiheng/TG-LLM.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI et al., 2023; Touvron et al., 2023b) have emerged with some reasoning capabilities  (Huang and Chang, 2023). However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems. In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Qin et al., 2021).
Figure 1: Our TG-LLM framework has two steps: text-to-temporal graph translation and temporal graph reasoning.
Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2023a). Similarly,  (Wei et al., 2023b) designs several specific types of prompts as prompt tuning.  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Given a reliable TG, the key challenges of teaching LLMs TR include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? Prior works  (Lewis et al., 2021) show that explicitly introducing knowledge in context enhances the performance of LLMs. In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., the temporal relation and time gap between the timestamps, the relative order of the gaps). (2) How can one teach LLM to perform deliberate reasoning? Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b). We adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. We further improve the model performance with graph data augmentation which mitigates the data deficiency in TR tasks.
@@
  • We design two approaches including Chain of Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
@@
Temporal Graph (sub; rel; obj; start/end; time):
@@
Table 1: Each sample from TGQA dataset is in the form of (story, temporal graph, questions, answers).
@@ 2 Dataset Construction
In this section, we present the construction pipeline for TGQA dataset that is fully controllable and requires minimal supervision for text-temporal graph alignment. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have more diverse categories and types of TR questions (Table 2), and the ground-truth timeline. We first split the large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), into subgraphs with a restriction on the number of events, and anonymize the entities to avoid data leakage. Then each story is translated from the subgraph by GPT-3.5  (Ouyang et al., 2022). By using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. Finally, to reduce the noise introduced from the misalignment between the subgraph and generated story, we propose a semi-automatic verification method.
@@
Step 1: Graph Splitting & Anonymization.  Existing temporal knowledge graphs (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) usually have a large size. To facilitate the learning process, we split YAGO11k into subgraphs for story generation. Specifically, given a certain entity, we find its neighbors within three hops, and extract all the events happening between them. Since LLMs are supposed to do reasoning instead of memorization, we ensure that no overlapping exists between the events in training, validation and test sets. Additionally, we notice the data leakage problem of LLMs, i.e., prior knowledge of the test data has been implicitly obtained from the extensive pre-training. Thus, an anonymization strategy, i.e., changing entity names into random ones of the same type, is adopted. We generate a global mapping for entity names with GPT-3.5. To avoid confusion, we adopt obscure names that do not exist in YAGO11k.
@@
Step 2: Graph-based Open QA Creation.  In TGQA, each sample is in the form of (temporal graph, story, questions, answers) (Table 1). Based on the given subgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the subgraph with the requirements to include all the events. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we separate the start and end time of the same event in prompt. On the other hand, to obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple question types (Table 2) with a rule-based Python script to generate the corresponding Qs and As.
@@
Step 3: Quality Control.  In TGQA, we might introduce noise from the misalignment between the given subgraph and generated story. To reduce the noise, we propose a semi-automatic verification method. Fully manual inspection is expensive, but by first asking LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM. Specifically, given a generated story, GPT-3.5 is queried on the time of each event in the graph. If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires manual verification. We proved the effectiveness of our semi-automatic pipeline with fully manual verification of the test stories. Note that supervision is only required for story-TG alignment verification, since all the QAs are generated from rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
@@ 3.1 Text-to-TG Translation
(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two ways: parsing with rules or using GPT-3.5. Rule-based parsing, if applicable, is more efficient and reliable, which is prioritized in our experiments.
@@
(3) TG Construction & Verification: We generate the TG using GPT-3.5 with ICL. Specifically, we provide several in-context demonstrations which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (Table 15). For those stories without explicit time expressions, we skip Step (2), and ask GPT-3.5 to provide the temporal order of events. Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking if GPT-3.5 fails.
Figure 2: In Chain of Thoughts (CoTs) bootstrapping, we accept CoTs leading to correct final answers and sample them according to their contrastive learning scores.
Supervised Fine-tuning. We first conducted experiments to decide the best in-context format of TG. It is found out that providing a chronological list of events helps LLM perform better TR. In addition, it offers advantages to separate the start and end time of the same event. With these steps, the TG in context is transferred into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023a) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
@@ 3.2.1 Bootstrapping Chain of Thoughts
Specifically, we first prepare ICL examples for high-quality CoT generation. To facilitate learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$, where $g$, $e$, $q$, $a$ denote TG, external knowledge, question and answer, respectively, and ask both GPT-3.5 and GPT-4 to provide diverse CoTs $\{c_{j}\}^{N_{i}}_{j=1}$. These CoTs will then be manually verified as ICL examples. Given a new training sample $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}})$, we bootstrap $K$ CoTs with final answers $\{(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})\}^{K}_{k=1}$ from GPT-3.5. We first refuse the CoTs leading to incorrect answers, i.e., $\hat{a}_{j^{\prime},k}\neq a^{*}_{j^{\prime}}$, where $a^{*}_{j^{\prime}}$ denotes the correct answer. For the accepted CoTs, we consider a weighted sampling strategy to balance usefulness and diversity. The sampling probability $P_{sample}(\cdot)$ is based on a contrastive learning $score$ (Eq. 1). The $score$ design, inspired by  (Wang et al., 2023), considers both perplexity $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq. 2). The definition of $G(\cdot)$ and $P(\cdot)$ are given in Eq. 3 and Eq. 4, respectively.
@@
$score(c_{k})=\log P\left(a^{*}\mid g,e,q,c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid g,e,q,c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid g,e,q,c_{k}\right)}$ (3)
$P\left(a\mid g,e,q,c_{k}\right)=\frac{1}{|a|}\underset{t_{l}\in a}{\sum}P\left(t_{l}\mid g,e,q,c_{k},t_{<l}\right)$ (4)
where $c_{k}$ denotes the current CoT in consideration; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
@@ 3.2.2 Graph Data Augmentation
Our framework involves two steps: text-to-TG translation and temporal knowledge graph reasoning. Notably, for the reasoning part, we train with ground-truth/verified TGs but infer on the estimated graphs. This discrepancy hurts the robustness of reasoning. Thus, we introduce some disturbances on the TGs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the LLM. We first investigate two types of disturbances: (i) remove irrelevant edges (Eq. 5) and (ii) replace edges by using relation synonyms (Eq. 6). An edge (event) is considered irrelevant if not involved in both QA and CoT.
@@
Figure 3: We have several strategies for graph data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
@@ 4.1 Experimental Setup
We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. Examples from other datasets are listed in Appendix B. We thoroughly evaluate our framework on all the datasets with a combination of the metrics, i.e., token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluation, i.e., selecting from a candidate set the final answer with the lowest perplexity as the prediction. The construction of the candidate set for all datasets and detailed explanations are listed in Appendix B.
We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al. (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models. We evaluated their few-shot ICL performance on the test set. To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the base model for the text-to-TG translation and temporal knowledge graph reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence. For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs. We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality. All the prompt templates are given in Appendix C.

### v4

@@ Abstract
While large language models (LLMs) have demonstrated remarkable reasoning capabilities, they are not without their flaws and inaccuracies. Recent studies have introduced various methods to mitigate these limitations. Temporal reasoning (TR), in particular, presents a significant challenge for LLMs due to its reliance on diverse temporal concepts and intricate temporal logic. In this paper, we propose TG-LLM, a novel framework towards language-based TR. Instead of reasoning over the original context, we adopt a latent representation, temporal graph (TG) that enhances the learning of TR. A synthetic dataset (TGQA), which is fully controllable and requires minimal supervision, is constructed for fine-tuning LLMs on this text-to-TG translation task. We confirmed in experiments that the capability of TG translation learned on our dataset can be transferred to other TR tasks and benchmarks. On top of that, we teach LLM to perform deliberate reasoning over the TGs via Chain-of-Thought (CoT) bootstrapping and graph data augmentation. We observed that those strategies, which maintain a balance between usefulness and diversity, bring more reliable CoTs and final results than the vanilla CoT distillation. [^1] Code and data are available at https://github.com/xiongsiheng/TG-LLM.
@@ 1 Introduction
As one of the fundamental abilities, temporal reasoning (TR) plays an important role in human perception. It is not just about understanding basic concepts such as ordering or duration; it extends to more intricate aspects, e.g., task planning or causal relation discovery. Recently, large language models (LLMs)  (Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a) have emerged with some reasoning capabilities  (Huang and Chang, 2023). However, there is observation that they still can not perform TR sufficiently well  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023), preventing their applications from solving complex real-world problems. In particular, TR requires a combination of various skills including mathematical and logical reasoning as well as commonsense knowledge  (Zhou et al., 2019; Vashishtha et al., 2020; Qin et al., 2021).
Figure 1: Our framework (TG-LLM) performs temporal reasoning in two steps: 1) Text-to-Temporal Graph translation: generate (relevant) temporal graph given the context and keyword (extracted from questions); 2) Temporal Graph Reasoning: perform Chain-of-Thought reasoning over the temporal graph.
Recent works mainly adopt general approaches to investigate and improve the TR capability of LLMs. For example,  (Wang and Zhao, 2023; Chu et al., 2023; Qiu et al., 2023) benchmark the leading LLMs on different TR tasks with standard Input/Output prompt, few-shot in-context learning (ICL) and Chain-of-Thought (CoT) reasoning  (Wei et al., 2022). Similarly,  (Wei et al., 2023) designs several specific types of prompts as prompt tuning.  (Li et al., 2023; Yuan et al., 2023) introduce pre-defined Python programs/rule-based templates to perform supervised fine-tuning (SFT). In addition,  (Tan et al., 2023a, b) adopt some extra strategies, which include specific pre-training, instruction tuning and reinforcement learning.
@@
Given a reliable TG, the key challenges of teaching TR to LLMs include: (1) How can one introduce the necessary arithmetic and commonsense knowledge involved in TR? Prior work  (Lewis et al., 2021) shows that explicitly introducing knowledge into context enhances the performance of LLMs. In this paper, we first identify all the valid time expressions, and then generate related knowledge (e.g., temporal relation and time gap between the timestamps, and the relative order of the gaps). (2) How can one teach LLM to perform deliberate reasoning? Generally, there exist two roadmaps: (i) translating natural language into logical statements, and using external symbolic engine for reasoning  Pan et al. (2023); (ii) using LLMs directly as the reasoning engine  (Zhu et al., 2023). For (i), the difficulty lies in accurate translation  Yang et al. (2023c) and the limited expressive power of formal logic. For (ii), there is no guarantee for the correctness of generated intermediate steps especially with insufficient training data  (Yang et al., 2023b). In this paper, we adopt (ii) with the proposed bootstrapping method to generate reliable intermediate steps for supervised fine-tuning. We further improve the model performance with graph data augmentation, which mitigates the data deficiency in TR tasks.
@@
  • We design two approaches including Chain-of-Thought bootstrapping and graph data augmentation to teach LLM to generate consistent and faithful CoTs, which brings better performance than the vanilla CoT distillation.
@@
Temporal Graph [sub; rel; obj; start/end; time]:
@@
Table 1: Each sample from TGQA dataset is in the form of (temporal graph, story, questions, answers).
@@ 2 Dataset Construction
In this section, we present the construction pipeline for TGQA dataset that is fully controllable and requires minimal supervision for text-temporal graph alignment. Compared with existing datasets  (Chen et al., 2021; Tan et al., 2023a), we have the ground-truth timelines and more diverse categories and types of TR questions (Table 2). More importantly, the pipeline can be used for various scenarios and tasks. We first split the large temporal knowledge graph, YAGO11k (Dasgupta et al., 2018), into subgraphs with a restriction on the number of events, and anonymize the entities to avoid data leakage. Then each story is translated from the subgraph by GPT-3.5  (Ouyang et al., 2022). By using some rule-based templates, we obtain reliable question and answer (QA) pairs from the graph. Finally, to reduce the noise introduced from the misalignment between the subgraph and generated story, we propose a semi-automatic verification method.
@@
Step 1: Graph Splitting & Anonymization.  Existing temporal knowledge graphs (Leetaru and Schrodt, 2013; Dasgupta et al., 2018; García-Durán et al., 2018) usually have a large size. To facilitate the learning process, we split YAGO11k into subgraphs for story generation. Specifically, given a certain entity, we find its neighbors within three hops, and extract all the events happening between them. Since we hope LLMs to do reasoning instead of memorization, it is ensured that no overlapping exists between the events in training, validation and test sets. Additionally, we notice the data leakage problem of LLMs, i.e., prior knowledge of the test data has been implicitly obtained from the extensive pre-training. Thus, an anonymization strategy, i.e., changing entity names into random ones of the same type, is adopted. For each relation, we generate a global mapping of entity names with GPT-3.5. To avoid confusion, we adopt obscure names that do not exist in YAGO11k.
@@
Step 2: Graph-based Open QA Creation.  In TGQA, each sample is in the form of (temporal graph, story, questions, answers) (Table 1). Based on the given subgraph, we generate a story and multiple QAs. We first ask GPT-3.5 to write a story based on the subgraph with the requirement to include all the events. It is observed that GPT-3.5 tends to ignore the end time of some events in the created story. To solve this problem, we separate the start and end time of the same event in the prompt. On the other hand, to obtain a comprehensive benchmark, we consider all types of temporal reasoning, which include sequencing, duration, frequency, simultaneity, temporal relation, comparative analysis and facts extraction. Given these categories, we design multiple question types (Table 2) with a rule-based Python script to generate the corresponding Qs and As.
@@
Step 3: Quality Control.  In TGQA, noise might be introduced from the misalignment between the given subgraph and LLM-generated story. To address this problem, we propose a semi-automatic verification method. Fully manual inspection is expensive, but by first utilizing LLM we can narrow down potential errors, and only manually inspect the set of unanswered questions by LLM. Specifically, given a generated story, GPT-3.5 is queried on the time of each event in the graph. If it cannot give the correct answer, we consider the event as possibly missing in the story, which requires further manual verification. We proved the effectiveness of our semi-automatic pipeline with fully manual verification of the test stories. Note that supervision is only required for story-TG alignment verification, since all the QAs are generated from rules. We show dataset statistics in Appendix A, and all the prompts involved in Appendix C.
@@ 3.1 Text-to-TG Translation
(1) Entity & Relation Extraction: To obtain the entity and relation in pre-defined QAs, we consider two strategies: parsing with rules or using GPT-3.5. Rule-based parsing, if applicable, is more efficient and reliable, which is prioritized in our experiments.
@@
(3) TG Construction & Verification: We generate the TG using GPT-3.5 with ICL. Specifically, we provide several in-context demonstrations which include a story and extracted entities, relations and time expressions as input and the corresponding TG as output (Table 15). For those stories without explicit time expressions, we skip Step (2), and ask GPT-3.5 to provide the temporal order of events. Similar to TGQA, we verify the generated TG by first asking GPT-3.5 the pre-defined QAs, and then manually checking the alignment if GPT-3.5 fails.
Figure 2: In Chain-of-Thought (CoT) bootstrapping, we only accept CoTs that lead to correct final answers and sample them according to their contrastive learning scores to balance usefulness and diversity.
Supervised Fine-tuning. We first conducted experiments to decide the best in-context format of TG. It is found out that providing a chronological list of events helps LLM perform better TR. In addition, it offers advantages to separate the start and end time of the same event. With these steps, the TG in context is transferred into a timeline with the alignment between entities, relations and times. We perform supervised fine-tuning (SFT) using Llama-2 model  (Touvron et al., 2023b) with Low-Rank Adaptation (LoRA)  (Hu et al., 2021). The input and output of the model are the story and aligned timeline, respectively. In experiments, we observe benefits from the SFT on our dataset to other temporal reasoning tasks.
@@ 3.2.1 Bootstrapping Chain of Thoughts
Specifically, we first prepare ICL examples for high-quality CoT generation. To facilitate the learning, the pre-defined QAs are classified into different categories $\{Q_{1},Q_{2},\cdots$, $Q_{M}\}$. For each category $Q_{i}$, we randomly choose $N_{i}$ samples $\{(g_{j},e_{j},q_{j},a_{j})\}^{N_{i}}_{j=1}$, where $g$, $e$, $q$, $a$ denote TG, external knowledge, question and answer, respectively, and ask both GPT-3.5 and GPT-4 to provide diverse CoTs $\{c_{j}\}^{N_{i}}_{j=1}$. These CoTs will then be manually verified as ICL examples. Given a new training sample $(g_{j^{\prime}},e_{j^{\prime}},q_{j^{\prime}})$, we bootstrap $K$ CoTs with final answers $\{(c_{j^{\prime},k},\hat{a}_{j^{\prime},k})\}^{K}_{k=1}$ from GPT-3.5. We first refuse the CoTs leading to incorrect answers, i.e., $\hat{a}_{j^{\prime},k}\neq a^{*}_{j^{\prime}}$, where $a^{*}_{j^{\prime}}$ denotes the correct answer. For the accepted CoTs, we consider a weighted sampling strategy to balance usefulness and diversity. The sampling probability $P_{sample}(\cdot)$ is based on a contrastive learning $score$ (Eq. 1). The $score$ design, inspired by  (Wang et al., 2023), considers both normalized probability $P(\cdot)$ of the correct answer and plausibility growth $G(\cdot)$ (Eq. 2). The definition of $G(\cdot)$ and $P(\cdot)$ are given in Eq. 3 and Eq. 4, respectively.
@@
$score(c_{k})=\log P\left(a^{*}\mid q^{\dagger},c_{k}\right)+\gamma G\left(c_{k}\right)$ (2)
$G\left(c_{k}\right)=\log\frac{P\left(a^{*}\mid q^{\dagger},c_{k}\right)}{\bar{P}_{\{a^{\prime}\in A^{\prime}\}}\left(a^{\prime}\mid q^{\dagger},c_{k}\right)}$ (3)
$\log P\left(a\mid q^{\dagger},c_{k}\right)=\frac{1}{|a|}\sum_{l}^{|a|}\log P\left(t_{l}\mid q^{\dagger},t_{<l}\right)$ (4)
where $c_{k}$ denotes the current CoT in consideration; $q^{\dagger}:=\{g,e,q\}$; $a^{\prime}$ denotes a certain wrong answer from the candidate set $A^{\prime}$; weight $\gamma$ is a hyperparameter; $t_{l}$ denotes the $l$-th token in $a$, and $t_{<l}$ denotes the sequence of tokens before $t_{l}$.
Figure 3: We further boost the model performance with several graph data augmentation strategies: remove irrelevant edges, use relation synonyms and change entities/times.
@@ 3.2.2 Graph Data Augmentation
Our framework involves two steps: text-to-TG translation and temporal graph reasoning. Notably, for the reasoning part, the model is trained over ground-truth/verified TGs but infers on the estimated graphs. This discrepancy actually hurts the robustness of reasoning. Thus, we introduce some disturbances on the TG during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the LLM. We first investigate two types of disturbances: (i) remove irrelevant edges (Eq. 5) and (ii) replace edges by using relation synonyms (Eq. 6). An edge (event) is considered irrelevant if not involved in both QA and CoT.
@@ 4.1 Experimental Setup
We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, and summaries. Examples from other datasets are listed in Appendix B. We thoroughly evaluate our framework on all the datasets with a combination of the metrics, i.e., token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluation, i.e., selecting from a candidate set the final answer with the lowest perplexity as the prediction. The rationale and detailed construction of the candidate set for all datasets are listed in Appendix B.
We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al. (2023b), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023). Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models. We evaluated their few-shot ICL performance on the test set. To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
@@
4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the base model for the text-to-TG translation and temporal graph reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence. For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs. We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality. All the prompt templates are given in Appendix C.